# Multiple merger coalescent inference of effective population size

### v1

@@ 3 Methods
Estimating $\alpha$ from tree topology only: By the end of Section 2, we showed that $\alpha$ affects both the tree shape and the distribution of coalescent times. While coalescent times are the sufficient statistics for inferring $N_{e}(t)$ in the standard coalescent process, coalescent times and tree topology (i.e. merging block sizes) are the sufficient statistics for estimating $N_{e}(t)$ and $\alpha$ under the Beta-coalescent. However, it is possible that in some applications, there is no reliable estimation of coalescent times and it may be preferable to estimate $\alpha$ from the tree topology only. In this case, we propose to estimate $\alpha$ using just the block sizes $\boldsymbol{m}$ by maximizing the following pseudo-likelihood under the Beta$(2-\alpha,\alpha)$-coalescent:
@@
Can we estimate $\alpha$ (or $\alpha$ and $N_{e}(t)$) from only coalescent times? The answer is no as the coalescent rate not only depends on coalescent times but on the total number of branches $A(t)$ obtained from $n$ and block sizes up to time $t$.
Hybrid estimation of $\alpha$ and $Ne(t)$: Our first approach for estimating $\alpha$ and $N_{e}(t)$ from a multiple merger genealogy is a hybrid approach, where we iteratively update $\alpha$ and $N_{e}(t)$. We first initialize $\alpha^{(0)}=\hat{\alpha}^{BS}$, to be the block size only MLE. Given $\alpha$, we then use INLA, as described in Section 2.2, and update $N_{e}(t)$ by the posterior median. Given the current value of $N_{e}(t)$, we then find $\alpha$ by maximum likelihood estimation, and iterate these last two steps until convergence.
Joint posterior inference of $N_{e}(t)$ and $\alpha$: We place a GMRF prior on $\log N_{e}(t)$ as described in Section 2.2 with a Gamma prior on precision parameter $\tau$, and a uniform $U(0,2)$ prior on $\alpha$. We approximate $\mathbb{P}(\alpha,N_{e}(t),\tau\mid\boldsymbol{g})$ via a Metropolis-within-Gibbs algorithm. The precision parameter and the GMRF are usually highly correlated leading Markov chain Monte Carlo methods to have poor convergence and slow mixing. To solve this problem, we sample from the conditional distribution of $\log N_{e}(t)$ and $\tau$ given all other model parameters and genealogy, with the split Hamiltonian Monte-Carlo (sHMC) method adapted from Lan et al. (2015). To sample $\alpha$ from the full marginal
@@
we discretize the parameter space of $\alpha\in[0,2]$ into intervals of length $0.005$ and approximate the denominator of Eq. 5 by the Riemann sum. That is, we set $\alpha_{m}$ to be the midpoint of interval $I_{m}=[0.005m,0.005(m+1)]$ for $m=1,...,400$. Then
@@ 4 Results
We test our methods on simulations and real applications. To evaluate their performance in estimating $\alpha$, we use the mean squared error (MSE). For evaluating the performance in estimating the effective population size, we compute four metrics: coverage, bias, deviance, and MSE at the grid points, with precise definitions in Appendix 7.2.
4.1 Simulations
To see how well we can estimate $\alpha$ from tree topology alone (block-size data), we generated 1000 realizations from the Beta-coalescent under different values of $\alpha,n$. Table 3 in Appendix 7.3 shows the mean, bias, and MSE of the estimates. Overall, we see $\hat{\alpha}^{BS}$ underestimates $\alpha$ and the accuracy of the estimator increases with $n$ and $\alpha$.
In the left plot in Figure 3, we show one isochronous tree with 50 tips, generated with $\alpha=1.5$ and exponential growth $N_{e}(t)=1000e^{-t}$. The block-size MLE is $\hat{\alpha}^{BS}\approx 1.615$ and the hybrid estimate is $\hat{\alpha}^{H}\approx 1.759$. The posterior median and mean of $\alpha$ inferred from MCMC are $1.5881$ and $1.5738$ respectively, which are the closest to the true $\alpha$ value. The right plot in Figure 3 depicts the reconstructed effective population size trajectories and their 95% credible regions estimated with each method, as well as the true trajectory. Table 2 in Appendix 7.3 details the performance measures calculated for inferred $N_{e}(t)$ with each method. We see that the inferring $N_{e}(t)$ using the true $\alpha$ has the lowest bias, deviance, and MSE, while out of the three proposed methods, MCMC performs the best and hybrid estimation the worst. This is expected since the $\hat{\alpha}^{H}$ is the worst estimate.
@@ 4.1 Simulations
Traj. BS MLE Hybrid MCMC
Unif 26.6% 32.3% 41.1%
Exp 30.6% 28.5% 40.9%
BB 32.7% 26.6% 40.7%
Table 1: Percentage of simulations where each method exhibits the lowest deviance for estimating $\alpha$.
Table 1 shows the proportion of simulations where each method (block size MLE, hybrid method, or MCMC) had the best performance according to deviance for estimating $\alpha$. We see that MCMC is the top performing method, with best performance in just over 40% of simulations. Block size MLE and hybrid estimation have similar performance. Table 4 in Appendix 7.3 shows the summary of performance statistics for estimating $N_{e}(t)$ (with the different sampling times combined). The value in boldface is the best performing (largest coverage, smallest deviance, and smallest MSE) out of the three methods. In terms of coverage, the MCMC method has the best performance, however all methods in general have good coverage. In terms of median deviance and mean MSE, the MCMC method is highly superior for the exponential and boom-and-boost trajectories, however we see the block-size MLE performs the best for the uniform trajectory, with the hybrid method not far behind.
@@ 5.1 Respiratory Syncytial Virus (RSV)
Here, we consider the genealogy shown in Figure 4(a) estimated from 266 RSV-A molecular sequences [^1] The data can be downloaded at https://nextstrain.org/rsv/a/G?f_country=Bangladesh,Cambodia,India,Laos,Myanmar,Nepal,Thailand,Vietnam sampled in South and Southeast Asia (mainly in India and Thailand) from 1991 to 2023. Sequences were aligned with MAFFT (Katoh et al., 2002) and the genealogy was reconstructed with IQ-Tree (Nguyen et al., 2015) paired with TreeTime (Sagulenko et al., 2018) for temporal inference. We accessed the webpage on April 9, 2024.. The block-size MLE is $\hat{\alpha}^{BS}=1.718$ and the hybrid estimate is $\hat{\alpha}^{H}=1.734$, and the MCMC posterior median and mean are 1.698 and 1.696. The inferred effective population size trajectories are shown in Figure 4(b). We see a rather steady increase until around 2010, with some oscillations indicative of the seasonality of RSV. The credible intervals obtained with MCMC are overall narrower than those obtained with the other two methods.
@@ 5.3 Japanese Sardine populations
We accessed the same mtDNA sequences from GenBank (accession nos. LC031518–LC031673), and reanalyzed the reproductive skew hypothesis with our methods. Out of the 156 total sequences, 106 were collected in 1990 and used in both of the previous papers for analysis, while the other 50 sequences, collected between 2010-2012 by the same authors were not included in the original study. We first used MAFFT to align the sequences of about 1200bp of length, then used IQ-Tree Version 1.6.12 and TreeTime to infer dated, multifurcating trees via maximum likelihood (Katoh et al., 2002; Nguyen et al., 2015; Sagulenko et al., 2018). We fix the mutation rate in TreeTime to be $2\times 10^{-7}$ per site per year. The related Indian oil sardine, Sardinella longiceps, estimated a mutation rate of $1\times 10^{-7}$ per site per year with a strict molecular clock (Sukumaran et al., 2016). Figure 6(a) and (b) show the inferred trees using 156 sequences and the sub-tree containing only sequences from 1990 respectively. The dated root is inferred to be around 29000 years ago, which roughly coincides around the time of the last glacial maximum.
@@ 6 Discussion
In this manuscript we propose three Bayesian methods to jointly estimate the effective population size and $\alpha$ parameter from a multifurcating tree under the Beta$(2-\alpha,\alpha)-$coalescent model. Two of the proposed approaches iterate estimation between $\alpha$ and $N_{e}(t)$, and the third approach approximates the posterior distribution of parameters via MCMC. The first two methods can be easily extended to any $\Lambda-$measure by computing the rates
@@
Finally, our methods assume that a genealogy is available without error. A natural future direction is to incorporate our approach into a full Bayesian approach that targets the posterior distribution of $N_{e}(t)$, $\alpha$, and $\mathbf{g}$ from molecular sequences directly. We anticipate this to be a challenging problem as the state space of multifurcating trees is much larger than the space of binary trees.

### v2

@@ 3 Methods
Estimating $\alpha$ from tree topology only: By the end of Section 2, we showed that $\alpha$ affects both the tree shape and the distribution of coalescent times. While coalescent times are the sufficient statistics for inferring $N_{e}(t)$ in the standard coalescent process, coalescent times and tree topology (i.e. merging block sizes) are the sufficient statistics for estimating $N_{e}(t)$ and $\alpha$ under the Beta-coalescent. However, it is possible that in some applications, there is no reliable estimation of coalescent times and it may be preferable to estimate $\alpha$ from the tree topology only. In this case, we propose to estimate $\alpha$ using just the block sizes $\boldsymbol{m}$ by maximizing the following pseudo-likelihood of the block sizes under the Beta$(2-\alpha,\alpha)$-coalescent (recall Equation 4):
@@
We know there is information about $\alpha$ in both the block sizes and the coalescent times. However, we show in Section 4, estimation of $\alpha$ based on tree topology only (block-sizes) is quite accurate.
Hybrid estimation of $\alpha$ and $N_{e}(t)$: Our first approach for estimating $\alpha$ and $N_{e}(t)$ from a multiple merger genealogy is a hybrid approach, where we iteratively update $\alpha$ and $N_{e}(t)$. We first initialize $\alpha^{(0)}=\hat{\alpha}^{BS}$, to be the block size only MLE. Given $\alpha$, we then use INLA, as described in Section 2.2, and update $N_{e}(t)$ by the posterior median. Given the current value of $N_{e}(t)$, we then find $\alpha$ by maximum likelihood estimation, and iterate these last two steps until convergence.
Joint posterior inference of $N_{e}(t)$ and $\alpha$: We place a GMRF prior on $\log N_{e}(t)$ as described in Section 2.2 with a Gamma prior on precision parameter $\tau$, and a uniform $U(0,2)$ prior on $\alpha$. We approximate $\mathbb{P}(\alpha,N_{e}(t),\tau\mid\boldsymbol{g})$ via a Metropolis-within-Gibbs algorithm, where $\boldsymbol{g}$ denotes the multifurcating genealogy. The precision parameter and the GMRF are usually highly correlated leading Markov chain Monte Carlo methods to have poor convergence and slow mixing. To solve this problem, we sample from the conditional distribution of $\log N_{e}(t)$ and $\tau$ given all other model parameters and genealogy, with the split Hamiltonian Monte-Carlo (sHMC) method adapted from Lan et al. (2015). To sample $\alpha$ from the full marginal
@@
we discretize the parameter space of $\alpha\in[0,2]$ into intervals of length $0.005$ and approximate the denominator of Eq. 7 by the Riemann sum. That is, we set $\alpha_{m}$ to be the midpoint of interval $I_{m}=[0.005m,0.005(m+1)]$ for $m=1,...,400$. The numerator of Equation 7 is proportional to the likelihood of Equation 3, and the proportionality constant cancels in the fraction. Then
@@ 4 Results
We first test our methods on simulated data. We implement a function that can simulate a Beta-coalescent genealogy under variable effective population size and heterochronous sampling. We use the true genealogy as input in simulation studies. To evaluate their performance in estimating $\alpha$, we use the mean squared error (MSE). For evaluating the performance in estimating the effective population size, we compute four metrics: coverage, bias, deviance, and MSE at the grid points, with precise definitions in Appendix 7.2.
To see how well we can estimate $\alpha$ from tree topology alone (block-size data), we generated 1000 realizations from the Beta-coalescent under different values of $\alpha,n$. Table 3 in Appendix 7.3 shows the mean, bias, and MSE of the estimated $\alpha$ using only the block-size data. Overall, we see $\hat{\alpha}^{BS}$ underestimates $\alpha$ and the accuracy of the estimator increases with $n$ and $\alpha$. When $\alpha=1.5$, the accuracy of this estimator is quite high for trees with 100 tips. When $\alpha=1.8$, the results are good for trees with 50 tips.
In the left plot in Figure 3, we show one isochronous tree with 50 tips, generated with $\alpha=1.5$ and exponential growth $N_{e}(t)=1000e^{-t}$. The block-size MLE is $\hat{\alpha}^{BS}\approx 1.615$ and the hybrid estimate is $\hat{\alpha}^{H}\approx 1.759$. The posterior median and mean of $\alpha$ inferred from MCMC are $1.594$ and $1.585$ respectively, which are the closest to the true $\alpha$ value. Figure 11 in Appendix 7.3 shows the posterior distribution of $\alpha$ from MCMC. The right plot in Figure 3 depicts the reconstructed effective population size trajectories and their 95% credible regions estimated with each method, as well as the true trajectory. Table 2 in Appendix 7.3 details the performance measures calculated for inferred $N_{e}(t)$ with each method. We see that the inferring $N_{e}(t)$ using the true $\alpha$ has the lowest bias, deviance, and MSE, while out of the three proposed methods, MCMC performs the best and hybrid estimation the worst. This is expected since the $\hat{\alpha}^{H}$ is the worst estimate. In Appendix 7.3 Figure 12, we show trace plots of $\alpha$ and $\log\;N_{e}(t)$ at a particular value of $t$ to demonstrate well-mixing.
@@
Coverage Deviance MSE
$N$ Method Unif. Exp. BB Unif. Exp. BB Unif. Exp. BB
$\alpha=1$ 20 BS MLE 0.92 0.96 0.95 0.4 1.23 1.57 19.34 117.14 287.95
Hybrid 0.91 0.84 0.84 0.55 3.76 5.03 31.7 1796.08 8197.65
MCMC 0.99 0.98 0.98 0.54 0.8 0.81 34.59 94.34 118.74
50 BS MLE 0.9 0.94 0.93 0.37 0.74 0.82 15.14 78.4 156.55
Hybrid 0.9 0.76 0.79 0.356 2.463 2.937 15.101 627.103 2389.601
MCMC 0.97 0.97 0.97 0.42 0.61 0.71 25.6 74 132.66
100 BS MLE 0.89 0.95 0.97 0.32 0.56 0.63 12.49 42.99 93.95
Hybrid 0.88 0.72 0.76 0.35 2.21 2.66 15.02 671.61 2454.44
MCMC 0.98 0.99 0.98 0.393 0.5 0.52 20.06 49.37 88.1
$\alpha=1.5$ 20 BS MLE 0.95 0.97 0.96 0.32 0.79 0.88 11.36 63.39 130.03
Hybrid 0.95 0.94 0.95 0.33 1.28 1.67 11.6 121.85 677.61
MCMC 0.99 0.97 0.98 0.33 0.54 0.57 14.22 63.44 85.45
50 BS MLE 0.95 0.97 0.96 0.26 0.66 0.68 7.47 60.97 125.58
Hybrid 0.95 0.94 0.95 0.27 1.05 1.12 8.13 115.94 497
MCMC 0.98 0.99 0.98 0.34 0.58 0.57 14.71 52.19 85.16
100 BS MLE 0.95 0.94 0.95 0.21 0.43 0.46 5.24 36.06 50.33
Hybrid 0.95 0.85 0.87 0.21 0.74 0.81 5.33 71.44 193.77
MCMC 0.99 0.98 0.98 0.28 0.41 0.41 9.84 34.78 49.26
$\alpha=1.8$ 20 BS MLE 0.98 0.98 0.98 0.23 0.66 0.74 6.44 60.77 153.67
Hybrid 0.99 0.97 0.98 0.24 0.74 0.86 6.3 56.85 207.42
MCMC 1 0.98 0.99 0.34 0.5 0.57 14.7 59.86 86.87
50 BS MLE 0.98 0.98 0.98 0.2 0.44 0.49 4.25 36.24 81.19
Hybrid 0.98 0.98 0.98 0.22 0.56 0.64 5 42.94 129.35
MCMC 0.99 0.99 0.99 0.29 0.41 0.41 10.25 27.56 56.11
100 BS MLE 0.97 0.96 0.96 0.17 0.43 0.43 3.6 23.18 37.24
Hybrid 0.97 0.96 0.97 0.16 0.48 0.47 3.27 25.03 52.15
MCMC 0.99 0.97 0.98 0.24 0.4 0.39 7.7 22.89 31.76
Table 1: Mean coverage, median deviance, and median MSE of estimated $N_{e}(t)$ for each trajectory and each method. The method with the best performance per $N,\alpha$ and trajectory combination is shown in boldface.
MCMC is the top performing method according to deviance for estimating $\alpha$ for over 40% of the simulations, with block size MLE and hybrid estimation have similar performances (see Table 4 in Appendix 7.3). Table 1 shows the summary of performance statistics for estimating $N_{e}(t)$ (with the different sampling times combined). The value in boldface is the best performing (largest coverage, smallest deviance, and smallest MSE) out of the three methods. In terms of coverage, the MCMC method has the best performance, however all methods in general have good coverage. In terms of median deviance and mean MSE, the MCMC method is highly superior for the exponential and boom-and-boost trajectories, however we see the block-size MLE performs the best for the uniform trajectory, with the hybrid method not far behind. One possible explanation is a lack of identifiability in the case of constant $N_{e}$: scaling $\alpha$ by $c$ and $N_{e}$ be $1/c$ would result in the same likelihood.
With regards to the computational cost of the methods, one run of the MCMC method (for 20,000 iterations) on a multifurcating tree with 100 tips takes approximately 15 minutes on one CPU node, while the BS MLE method takes less than 10 seconds, and the hybrid method takes 3 minutes. Computational complexity of all methods is roughly linear in $n$, the number of tips.
@@ 5.1 Respiratory Syncytial Virus (RSV)
Here, we consider the genealogy shown in Figure 4(a) estimated from 266 RSV-A molecular sequences [^1] The data can be downloaded at https://nextstrain.org/rsv/a/G?f_country=Bangladesh,Cambodia,India,Laos,Myanmar,Nepal,Thailand,Vietnam sampled in South and Southeast Asia (mainly in India and Thailand) from 1991 to 2023. Sequences were aligned with MAFFT (Katoh et al., 2002) and the genealogy was reconstructed with IQ-Tree (Nguyen et al., 2015). Finally, we used TreeTime (Sagulenko et al., 2018) for temporal inference and branches of length zero were collapsed. We accessed the webpage on April 9, 2024. The block-size MLE is $\hat{\alpha}^{BS}=1.718$ and the hybrid estimate is $\hat{\alpha}^{H}=1.734$, and the MCMC posterior median and mean are 1.698 and 1.696. The inferred effective population size trajectories are shown in Figure 4(b). We see a rather steady increase until around 2010, with some oscillations indicative of the seasonality of RSV. The credible intervals obtained with MCMC are overall narrower than those obtained with the other two methods.
@@ 5.3 Japanese Sardine populations
We accessed the same mtDNA sequences from GenBank (accession nos. LC031518–LC031673), and reanalyzed the reproductive skew hypothesis with our methods. Out of the 156 total sequences, 106 were collected in 1990 and used in both of the previous papers for analysis, while the other 50 sequences, collected between 2010-2012 by the same authors were not included in the original study. We first used MAFFT to align the sequences of about 1200bp of length, then used IQ-Tree Version 1.6.12 and TreeTime to infer dated, multifurcating trees via maximum likelihood (Katoh et al., 2002; Nguyen et al., 2015; Sagulenko et al., 2018). Again, branches of length zero were collapsed. We fix the mutation rate in TreeTime to be $2\times 10^{-7}$ per site per year. The related Indian oil sardine, Sardinella longiceps, estimated a mutation rate of $1\times 10^{-7}$ per site per year with a strict molecular clock (Sukumaran et al., 2016). Figure 6(a) and (b) show the inferred trees using 156 sequences and the sub-tree containing only sequences from 1990 respectively. The dated root is inferred to be around 29000 years ago, which roughly coincides around the time of the last glacial maximum.
@@ 6 Discussion
In this manuscript we propose three Bayesian nonparametric methods to jointly estimate the effective population size and $\alpha$ parameter from a multifurcating tree under the Beta$(2-\alpha,\alpha)-$coalescent model. Two of the proposed approaches iterate estimation between $\alpha$ and $N_{e}(t)$, and the third approach approximates the posterior distribution of parameters via MCMC. The first two methods can be easily extended to any $\Lambda$-measure by computing the rates
@@
Finally, our methods assume that a genealogy is available without error, or rather is presented to the practitioner. We include a small simulation example in Appendix 7.4 comparing parameter estimates from the true simulated genealogy versus reconstructed genealogies from simulated molecular sequences. As expected, there is larger uncertainty when genealogies are estimated, in particular in the estimation of $N_{e}(t)$. A natural future direction is to incorporate our approach into a full Bayesian approach that targets the posterior distribution of $N_{e}(t)$, $\alpha$, and $\mathbf{g}$ from molecular sequences directly. We anticipate this to be a challenging problem as the state space of multifurcating trees is much larger than the space of binary trees.